Please be very detailed when explaining this I want to understand to the best of my ability.Q 18.11: The energy given off by 300 g of an alloy as it cools by 50°C raises the temperature of 300 g of water from 30°C to
Answer #1

Given,

The mass, m1 = 300 g = 0.3 kg

m2 = 300 g = 0.3 kg

\DeltaT1 = 50 0C

\DeltaT2 = 40 - 30 = 10 0C

The specific heat of the alloy is,

Q = mc\DeltaT

So,

m1c1\DeltaT1 = m2c2\DeltaT2

300 * c1 * 50 = 300 * 1 * 10

c1 = 0.2 cal / g0C

Answer: B) 0.2 cal / g0C